OpenStudy (loser66):

How are they different? Find inf, sup of S 1) S= {1/n | n in N} 2) S ={1/n | n in Z, n >0} 3) S ={1/n | n in Z , n> = 0} Please, help

OpenStudy (mathmath333):

sry 1.) inf =0,sup=1 ???

OpenStudy (mathmath333):

inf means lowest?

OpenStudy (mathmath333):

and vice versa

OpenStudy (loser66):

inf means the greatest lower bound

OpenStudy (mathmath333):

so m i correct

OpenStudy (loser66):

Yes, but put it in logic, please

OpenStudy (mathmath333):

i never heard of this questions

OpenStudy (mathmath333):

well S will be maximum only at n=1

OpenStudy (loser66):

Like: Claim inf S = 0, Proof: \(\forall n \in N\) \(\dfrac{1}{n}<0 \)

OpenStudy (mathmath333):

and s will be lowest on 1/infinty =0

OpenStudy (loser66):

but sup S is not 1, I think

OpenStudy (loser66):

if I take n = 0.5 , n is in N, and 1/n = 2 > 1, so that 1 can't be upper bound of S

OpenStudy (mathmath333):

nut u specified n is in set of natural numbers and 0.5 is not natural number

OpenStudy (loser66):

oh yes, I just realize that, I am sorry, my bad

OpenStudy (mathmath333):

i mean but u specified ,lol

OpenStudy (loser66):

Ok, next n in Z

OpenStudy (loser66):

The same, right?

OpenStudy (loser66):

for the third case, \(n\in Z, n\geq 0\) is the same, still. right?

OpenStudy (mathmath333):

yes 1 and 2 are same

OpenStudy (mathmath333):

no

OpenStudy (mathmath333):

i think its inf=1 and sup =1/0=infinity

OpenStudy (mathmath333):

not sure lol

OpenStudy (loser66):

Ok, thank you so much. I got what happen.

OpenStudy (mathmath333):

http://en.wikipedia.org/wiki/Infimum_and_supremum this helped me lol

OpenStudy (mathmath333):

hey wait

OpenStudy (mathmath333):

for third it will be inf =1/infinity=0 and sup=1/0=infinity

OpenStudy (mathmath333):

tricky

OpenStudy (loser66):

when n goes to 0 , 1/n either goes to \(\infty\) or \(-\infty\) so that sup, inf don't exists.

OpenStudy (loser66):

but if we break it down to 2 intervals \([-\infty, 0]\) and \([0,\infty]\) then we have inf for each = 0

OpenStudy (mathmath333):

for third question but n>=0 so it will be positive infinity right?

OpenStudy (loser66):

yes,

OpenStudy (mathmath333):

so it means infium exists for question 3 as zero

OpenStudy (loser66):

yes,

OpenStudy (loser66):

hey, I am the Asker, not you. hehehe... You are helping me. How can it turn to I answer your question situation?? hahahha

OpenStudy (mathmath333):

lol its typical way of answering/interacting...dont like other boring ways

OpenStudy (loser66):

Anyway!!! thanks for your help.

OpenStudy (mathmath333):

welcome your